David owns $20,000 worth of 10-year bonds of Ace Corporation. These bonds pay interest every 6 months at the rate of 2%/year (simple interest).
How much income will David receive from this investment every 6 months?


How much interest will David receive over the life of the bonds?

Answer by ikleyn(52915)

.
David owns $20,000 worth of 10-year bonds of Ace Corporation.
These bonds pay interest every 6 months at the rate of 2%/year (simple interest).

(a) How much income will David receive from this investment every 6 months?

(b) How much interest will David receive over the life of the bonds?
~~~~~~~~~~~~~~


(a)  Use the formula for simple interest  I = P*r*t.


     In case (a),  I = 20000*0.02*0.5 = 200.


     David receives $200 simple interest every 6 months (0.5 of an year).    ANSWER




(b)  To answer (b), multiply $200 by 2 and by 10:   2 * 10 * $200 = $4000  simple interest in 10 years.    ANSWER


Solved.

------------------

To see many other similar  (and different)  solved problems on simple interest accounts,  look into the lesson
    - Simple interest percentage problems
in this site.

Learn the subject from there.